Find the coordinates of the points on the curve 𝑦 = 3π‘₯Β³ βˆ’ 2π‘₯Β² βˆ’ 12π‘₯ + 2 where



the normal is parallel to the line 𝑦 =βˆ’ π‘₯ + 1.

1
Expert's answer
2022-03-11T05:48:11-0500

Find the coordinates of the points on the curve 𝑦 = 3π‘₯Β³ βˆ’ 2π‘₯Β² βˆ’ 12π‘₯ + 2 where

the normal is parallel to the line 𝑦 =βˆ’ π‘₯ + 1.


The derivative

"y'=9x^2-4x-12"

The slope of the normal is -1, hence the slope of the tangent line is "-\frac{1}{-1}=1" .

Then

"y'=-1\Rightarrow 9x^2-4x-12=1\Rightarrow 9x^2-4x-13=0\Rightarrow
x\in \{-1,\frac{13}{9}\}"

The points are

"(-1,9), (\frac{13}{9},-\frac{2543}{243})"